# Resonant leptogenesis in minimal $U(1)_{X}$ extensions of the Standard Model

### v1

@@ Abstract
We investigate a general $U(1)_{X}$ scenario where we introduce three generations of Standard Model (SM) singlet Right Handed Neutrinos (RHNs) to generate the light neutrino mass through the seesaw mechanism after the breaking of $U(1)_{X}$ and electroweak symmetries. In addition to that, a general $U(1)_{X}$ scenario involves an SM-singlet scalar field and due to the $U(1)_{X}$ symmetry breaking the mass of a neutral beyond the SM (BSM) gauge boson $Z^{\prime}$ is evolved. The RHNs, being charged under $U(1)_{X}$ scenario, can explain the origin of observed baryon asymmetry through the resonant leptogenesis process. Applying observed neutrino oscillation data we study $Z^{\prime}$ and BSM scalar induced processes to reproduce the observed baryon asymmetry. Hence we estimate bounds on the $U(1)_{X}$ gauge coupling $(g_{X})$ and the mass of the $Z^{\prime}$ $(M_{Z^{\prime}})$ for different $U(1)_{X}$ charges and benchmark masses of RHN and SM-singlet scalar. Finally we compare our results with limits obtained from the existing limits from LEP-II and LHC. We find that depending on the $U(1)_{X}$ charges, the masses of RHNs and SM-singlet scalar resonant leptogenesis could provide stronger limit on $g_{X}$ for $M_{Z^{\prime}}>5.8$ TeV which could be probed by high energy scattering experiment in future.
@@ I Introduction
The origin of baryon asymmetry of the universe is an unsolved puzzle of modern cosmology. The WIMP satellite WMAP:2010qai observed that the ratio of the baryon minus anti-baryon density $(n_{B-\bar{B}})$ over the entropy density of the universe $(s)$ to be $Y_{B}=8.7\times 10^{-11}$ at an accuracy of $10\%$ precision level. A strong first order phase transition could be required to explain the origin of baryon (B) asymmetry induced by the electroweak baryogenesis Cohen:1993nk ; Funakubo:1996dw ; Trodden:1998ym within the Standard Model (SM) framework, however, the observation of the SM Higgs mass around 125 GeV CMS:2012qbp ; ATLAS:2012yve does not indicate such a phenomenon. As a result electroweak baryogenesis is possibly ruled out under the SM scenario.
@@
In addition to the anomaly associated with the observed baryon asymmetry there are interesting developments in the lepton sector of the SM which are observed in the form of tiny neutrino mass, neutrino flavor oscillation and mixing ParticleDataGroup:2022pth and they can not be explained within the SM framework. As a result an approach of stepping beyond the SM (BSM) was important which evolved in the form of seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia ; Glashow:1979nm where the origin of tiny neutrino mass could be explained through the suppression of a heavy Majorana mass scale. In this case the SM is extended by SM-singlet heavy Majorana type Right Handed Neutrinos (RHNs) at the intermediate scale which create lepton number violation by unit 2 through a dimension five Weinberg operator Weinberg:1979sa .
@@
An attempt to connect the baryon asymmetry of the universe and possibility of tiny neutrino mass could be established by the seesaw scenario by an attractive possibility called leptogenesis Fukugita:1986hr . The out of equilibrium decay of the Majorana type RHNs can generate lepton (L) asymmetry in the universe which can be converted into baryon asymmetry through the sphaleron transition violating $(B+L)$ quantum number Manton:1983nd ; Klinkhamer:1984di . Hence the ratio between baryon and lepton conversion rate can be $-\frac{28}{79}$ following Khlebnikov:1988sr ; Iso:2010mv ; Chauhan:2021xus for three (one) generations of SM fermions (Higgs). In this context we mention that thermal leptogenesis imparts a lower bound on the Majorana neutrinos $(M_{N})$ to generate sufficient amount of baryon asymmetry as $M_{N}>10^{9}$ GeV Plumacher:1996kc ; Buchmuller:2002rq ; Buchmuller:2004nz . Such heavy Majorana neutrinos can not be observed at the hadron and lepton colliders.
@@
In order to test heavy Majorana neutrinos at high energy colliders we need to think possible alternative scenarios which may explain the origin of small neutrino mass and observed baryon asymmetry simultaneously. General U$(1)$ extensions of the SM are such scenarios where these aspects can be addressed nicely. In this case, the SM is extended by a general $U(1)_{X}$ gauge group where three generations of the RHNs are introduced with an SM-singlet BSM scalar. These RHNs not only help generating the tiny neutrino mass though the seesaw mechanism after the general $U(1)$ symmetry breaking, but also cancel the gauge gravity anomalies Oda:2015gna ; Das:2016zue . If general $U(1)$ symmetry breaking scale lies around the TeV scale, the corresponding neutral gauge boson commonly known as $Z^{\prime}$ and the RHNs can be produced at high energy colliders Das:2022rbl . In this scenario there is another interesting aspect which affects the $Z^{\prime}$ searches. After the anomaly cancellation we find that the left and right handed fermions of the model interact differently with the $Z^{\prime}$ boson. The general $U(1)$ scenarios considered in this article are at the TeV scale prohibiting thermal leptogenesis to occur because in this case the RHNs are lighter than the bound of $10^{10}$ GeV. The CP asymmetry parameter in this case is proportional to the modulus squared of the Dirac Yukawa coupling $(Y_{D})$ between the RHNs and the SM lepton doublet. Due to the smallness of $Y_{D}$, the right amount of baryon asymmetry of the universe can not be generated. It has been also found that if two RHNs are almost degenerate in masses, an enhancement Flanz:1996fb of the CP asymmetry parameter can take place making the leptogenesis scenario viable for the TeV scale RHNs. This is called resonant leptogenesis Pilaftsis:1997jf ; Pilaftsis:2003gt . In this case maximum attainable enhancement could be achieved if mass difference between the two generations of the RHNs is $(\mathcal{O}(\Gamma_{N_{i}}))$, the total decay width $(\Gamma_{N_{i}})$ of either of the generations of the RHNs $(N_{i})$. Hence in principle by tuning the mass difference between any two generations of the RHNs, the CP asymmetry could be attained around $\mathcal{O}(1)$, however, in the presence of general $U(1)$ extensions the lepton asymmetry through RHN decay is suppressed due to the interaction with the TeV scale $Z^{\prime}$ gauge boson. As a result such a scenario makes the generation of the CP asymmetry of the universe non-trivial Blanchet:2009bu .
@@
In this paper we study resonant leptogenesis in general $U(1)$ extensions of the SM where $U(1)_{X}$ and $U(1)_{q+xu}$ symmetry breaks at the TeV scale. Solving Boltzmann equations and considering different model parameters we show the dependance of baryon asymmetry via leptogenesis on neutrino Dirac Yukawa couplings and TeV scale nearly degenerate heavy Majorana neutrino masses. In this analysis we involve neutrino oscillation parameters in order to relate the baryon asymmetry via leptogenesis with realistic data. In this analysis we also study the role of the Dirac CP phase in the $U_{\rm{PMNS}}$ matrix in order to explain baryon asymmetry of the universe.
@@
We organize the paper in the following way. In Sec. II we describe the model. In Sec. III we analyse the resonant leptogenesis scenario. In Sec. IV we apply neutrino oscillation data to investigate the role of Dirac CP phase to generate right amount of baryon asymmetry of the universe. We discuss our results in Sec. IV and conclude it in Sec. V.
@@ II General $U(1)$ extensions of Standard Model
respectively. Solving Eqs. (2) and (3) we express the fermionic $U(1)_{X}$ charges in terms of $x_{H}$ and $x_{\Phi}$ and the $U(1)_{X}$ charges of the SM fermions can be written as linear combination of the $U(1)_{Y}$ and B--L charges. Hence we find that left and right handed fermions are differently charged under the $U(1)_{X}$ gauge group and as a result they interact differently with the associated neutral BSM gauge boson $Z^{\prime}$ manifesting a chiral nature of the model. Such a chiral behavior affect any interaction process mediated by the $Z^{\prime}$ associating the fermions of this model. In this scenario, if we fix $x_{\Phi}=1$ and $x_{H}=0$ we can reproduce the B--L scenario and with $x_{H}=-2$ we find that the $U(1)_{X}$ charges of the left handed fermions become zero, however, the right handed fermions have nonzero $U(1)_{X}$ charges. This is called $U(1)_{\rm R}$ scenario. In addition to that there are other properties of this model if we change $x_{H}$ keeping $x_{\Phi}=1$ fixed as 1. For example, in case of $x_{H}=-1$, the $U(1)_{X}$ charge of $e_{R}$ vanishes, in case of $x_{H}=-0.5$, the U$(1)_{X}$ charge of $u_{R}$ vanishes and in case of $x_{H}=1$, the $U(1)_{X}$ charge of $d_{R}$ vanishes, respectively. However, for these respective $x_{H}$ values remaining fermions interact with the $Z^{\prime}$. The $U(1)_{X}$ gauge coupling is considered as $g_{X}$ for this model. This is a free parameter and it appears as either $g^{\prime}x_{H}$ or $g^{\prime}x_{\Phi}$ in the kinetic part of the Lagrangian. The renormalizable scalar potential in this scenario can be written as
@@
where $q_{f_{L}}$ is the $U(1)_{X}$ charge of the SM lepton doublet, $Z^{\prime}$ gauge boson can decay into a pair of heavy Majorana neutrinos and the partical decay width is given by
@@ IV Results and discussions
$\frac{g^{\prime 2}}{{M_{Z^{\prime}}}^{2}-s}[\overline{e}\gamma^{\mu}(\tilde{x}_{\ell}P_{L}+\tilde{x}_{e}P_{R})e][\overline{f}\gamma_{\mu}(\tilde{x}_{f_{L}}P_{L}+\tilde{x}_{f_{R}}P_{R})f]\,,$ (56)
@@
$M_{Z^{\prime}}^{2}\ \gtrsim\ \frac{{g^{\prime}}^{2}}{4\pi}|{x_{e_{A}}}x_{f_{B}}|(\Lambda_{AB}^{f\pm})^{2}\,,$ (57)
@@
We calculate limits on $g_{X}-M_{Z^{\prime}}$ in the $U(1)_{X}$ scenario for different $x_{H}$ from the dilepton and dijet searches in ATLAS and CMS ATLAS:2019erb ; CMS:2019tbu experiments of the the LHC. The corresponding cross sections for these scenarios are estimated as $\sigma_{\rm Model}$ involving the $Z^{\prime}$ contribution for different $x_{H}$ from $U(1)_{X}$ model considering a trial value of the general $U(1)_{X}$ gauge coupling $g_{\rm{Model}}$ at $\sqrt{s}=14$ TeV varying $M_{Z^{\prime}}$. Comparing these estimated cross sections with the observed cross sections from the LHC $(\sigma_{\rm{Obs.}})$ to estimate limits on the gauge coupling for different $M_{Z^{\prime}}$ and $x_{H}$ following
$g^{\prime}\ =\ \sqrt{g_{\rm{Model}}^{2}\left(\frac{\sigma_{\rm{Obs.}}}{\sigma_{\rm{Model}}}\right)}.$ (58)
to estimate $95\%$ constraints on the $g_{X}-M_{Z^{\prime}}$. The LHC limits for the dilepton (dotted) and dijet (dot-dashed) events from CMS (magenta) and ATLAS (blue) are shown in Fig. 2. The dilepton bounds provide strong constrains on $g_{X}-M_{Z^{\prime}}$ plane for different $x_{H}$.
@@
Finally from Fig. 2 we compare the limits obtained from resonant leptogenesis and collider searches for different $x_{H}$. We find that limits obtained from resonant leptogenesis become stronger than the limits obtained from LEP-II, LHC for $M_{Z^{\prime}}>5.8$ TeV for $x_{H}=-2$ and $-1$. The constraints obtained from LEP-II become stronger than the CMS limits for $x_{H}\leq 0$, ruling out some parts of the resonant leptogenesis parameters for $x_{H}=0$ and $1$, respectively. We find that For $x_{H}=2$, LEP-II limits rule out the limits obtained from the resonant leptogenesis. We find that prospective limits in case of 250 GeV ILC can be comparable with the limits obtained by the resonant lespogenesis scenario. Hence these limits could be probed at 250 GeV ILC for $x_{H}=-1$. Other prospective limits from ILC at different center of mass energies more than 250 GeV and different $x_{H}$ could provide stronger constraints than the limits obtained from resonant leptogenesis.

### v2

@@ Abstract
We investigate a general $U(1)_{X}$ scenario where we introduce three generations of Standard Model (SM) singlet Right Handed Neutrinos (RHNs) to generate the light neutrino mass through the seesaw mechanism after the breaking of $U(1)_{X}$ and electroweak symmetries. In addition to that, a general $U(1)_{X}$ scenario involves an SM-singlet scalar field and due to the $U(1)_{X}$ symmetry breaking the mass of a neutral beyond the SM (BSM) gauge boson $Z^{\prime}$ is evolved. The RHNs, charged under the $U(1)_{X}$ scenario, can explain the origin of observed baryon asymmetry through the resonant leptogenesis process. Applying observed neutrino oscillation data we study $Z^{\prime}$ and BSM scalar-induced processes to reproduce the observed baryon asymmetry. Hence we estimate bounds on the $U(1)_{X}$ gauge coupling $(g_{X})$ and the mass of the $Z^{\prime}$ $(M_{Z^{\prime}})$ for different $U(1)_{X}$ charges and benchmark masses of the RHN and SM-singlet scalar. Finally, we compare our results with limits obtained from the existing limits from LEP-II and LHC. We find that depending on the $U(1)_{X}$ charges, the masses of RHNs and SM-singlet scalar resonant leptogenesis could provide a stronger limit on $g_{X}$ for $M_{Z^{\prime}}>5.8$ TeV which could be probed by high energy scattering experiment in future.
@@ I Introduction
The origin of baryon asymmetry of the universe is an unsolved puzzle of modern cosmology. The WIMP satellite WMAP:2010qai observed that the ratio of the baryon minus anti-baryon density $(n_{B-\bar{B}})$ over the entropy density of the universe $(s)$ is $Y_{B}=8.7\times 10^{-11}$ at an accuracy of $10\%$ precision level. A strong first-order phase transition could be required to explain the origin of baryon (B) asymmetry induced by the electroweak baryogenesis Cohen:1993nk ; Funakubo:1996dw ; Trodden:1998ym within the Standard Model (SM) framework, however, the observation of the SM Higgs mass around 125 GeV CMS:2012qbp ; ATLAS:2012yve does not indicate such a phenomenon. As a result, electroweak baryogenesis is possibly ruled out in the SM scenario.
@@
In addition to the anomaly associated with the observed baryon asymmetry there are interesting developments in the lepton sector of the SM which are observed in the form of tiny neutrino mass, neutrino flavor oscillation and mixing ParticleDataGroup:2022pth and can not be explained within the SM framework. As a result an approach of stepping beyond the SM (BSM) was important which evolved in the form of seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia ; Glashow:1979nm where the origin of tiny neutrino mass could be explained through the suppression of a heavy Majorana mass scale. In this case the SM is extended by SM-singlet heavy Majorana type Right Handed Neutrinos (RHNs) at the intermediate scale which create lepton number violation by unit 2 through a dimension five Weinberg operator Weinberg:1979sa .
@@
An attempt to connect the baryon asymmetry of the universe and possibility of tiny neutrino mass could be established by the seesaw scenario by an attractive possibility called leptogenesis Fukugita:1986hr . The out of equilibrium decay of the Majorana type RHNs can generate lepton (L) asymmetry in the universe which can be converted into baryon asymmetry through the sphaleron transition violating $(B+L)$ quantum number Manton:1983nd ; Klinkhamer:1984di . Hence, the ratio between the baryon and lepton conversion rate can be $-\frac{28}{79}$ following Khlebnikov:1988sr ; Iso:2010mv ; Chauhan:2021xus for three (one) generations of SM fermions (Higgs). In this context, we mention that thermal leptogenesis imparts a lower bound on the Majorana neutrinos $(M_{N})$ to generate sufficient amount of baryon asymmetry as $M_{N}>10^{9}$ GeV Plumacher:1996kc ; Buchmuller:2002rq ; Buchmuller:2004nz . Such heavy Majorana neutrinos can not be observed at the hadron and lepton colliders.
@@
In order to test heavy Majorana neutrinos at high energy colliders we need to think possible alternative scenarios which may explain the origin of small neutrino mass and observed baryon asymmetry simultaneously. General U$(1)$ extensions of the SM are such scenarios in which these aspects can be addressed nicely. In this case, the SM is extended by a general $U(1)_{X}$ gauge group where three generations of the RHNs are introduced with an SM-singlet BSM scalar. These RHNs not only help to generate the tiny neutrino mass though the seesaw mechanism after the general $U(1)$ symmetry breaking, but also cancel the gauge gravity anomalies Oda:2015gna ; Das:2016zue . If general $U(1)$ symmetry breaking scale lies around the TeV scale, the corresponding neutral gauge boson commonly known as $Z^{\prime}$ and the RHNs production at high energy colliders. We extensively studied the RHN production form $Z^{\prime}$ at different hadron and lepton colliders in Das:2022rbl applying detector simulations which are not taken into account in a recent article Chauhan:2024jfq while following the flow-chart of Das:2022rbl with inadequate insights on the results. In this scenario there is another interesting aspect which affects the $Z^{\prime}$ searches. After the anomaly cancellation we find that the left and right handed fermions of the model interact differently with the $Z^{\prime}$ boson. The general $U(1)$ scenarios considered in this article are at the TeV scale prohibiting thermal leptogenesis to occur, because in this case the RHNs are lighter than the bound of $10^{10}$ GeV. The CP asymmetry parameter in this case is proportional to the modulus squared of the Dirac Yukawa coupling $(Y_{D})$ between the RHNs and the SM lepton doublet. Due to the smallness of $Y_{D}$, the right amount of baryon asymmetry of the universe can not be generated. It has also been found that if two RHNs are almost degenerate in masses, an enhancement Flanz:1996fb of the CP asymmetry parameter can take place making the leptogenesis scenario viable for the TeV scale RHNs. This is called resonant leptogenesis Pilaftsis:1997jf ; Pilaftsis:2003gt . In this case maximum attainable enhancement could be achieved if mass difference between the two generations of the RHNs is $(\mathcal{O}(\Gamma_{N_{i}}))$, the total decay width $(\Gamma_{N_{i}})$ of either of the generations of the RHNs $(N_{i})$. Hence in principle by tuning the mass difference between any two generations of the RHNs, the CP asymmetry could be attained around $\mathcal{O}(1)$, however, in the presence of general $U(1)$ extensions the lepton asymmetry through RHN decay is suppressed due to the interaction with the TeV scale $Z^{\prime}$ gauge boson. As a result such a scenario makes the generation of the CP asymmetry of the universe non-trivial Blanchet:2009bu .
@@
In this paper, we study resonant leptogenesis in general $U(1)$ extensions of the SM where $U(1)_{X}$ and $U(1)_{q+xu}$ symmetry breaks at the TeV scale. Solving Boltzmann equations and considering different model parameters we show the dependence of baryon asymmetry via leptogenesis on neutrino Dirac Yukawa couplings and TeV scale nearly degenerate heavy Majorana neutrino masses. In this analysis we involve neutrino oscillation parameters in order to relate the baryon asymmetry via leptogenesis with realistic data. In this analysis we also study the role of the Dirac CP phase in the $U_{\rm{PMNS}}$ matrix in order to explain baryon asymmetry of the universe.
@@
We organize the paper in the following way. In Sec. II we describe the model. In Sec. III we analyse the resonant leptogenesis scenario. In Sec. IV we apply neutrino oscillation data to investigate the role of Dirac CP phase to generate the right amount of baryon asymmetry of the universe. We discuss our results in Sec. IV and conclude it in Sec. V.
@@ II General $U(1)$ extensions of Standard Model
respectively. Solving Eqs. (2) and (3) we express the fermionic $U(1)_{X}$ charges in terms of $x_{H}$ and $x_{\Phi}$ and the $U(1)_{X}$ charges of the SM fermions can be written as linear combination of the $U(1)_{Y}$ and B--L charges. Hence we find that left and right handed fermions are differently charged under the $U(1)_{X}$ gauge group and as a result they interact differently with the associated neutral BSM gauge boson $Z^{\prime}$ manifesting a chiral nature of the model. Such a chiral behavior affect any interaction process mediated by the $Z^{\prime}$ associating the fermions of this model. In this scenario, if we fix $x_{\Phi}=1$ and $x_{H}=0$ we can reproduce the B--L scenario and with $x_{H}=-2$ we find that the $U(1)_{X}$ charges of the left handed fermions become zero, however, the right handed fermions have nonzero $U(1)_{X}$ charges. This is called $U(1)_{\rm R}$ scenario. In addition to that there are other properties of this model if we change $x_{H}$ keeping $x_{\Phi}=1$ fixed as 1. For example, in case of $x_{H}=-1$, the $U(1)_{X}$ charge of $e_{R}$ vanishes, in case of $x_{H}=-0.5$, the U$(1)_{X}$ charge of $u_{R}$ vanishes and in case of $x_{H}=1$, the $U(1)_{X}$ charge of $d_{R}$ vanishes, respectively. However, for these respective $x_{H}$ values remaining fermions interact with the $Z^{\prime}$. The $U(1)_{X}$ gauge coupling is considered as $g_{X}$ for this model. This is a free parameter and it appears as either $g^{\prime}x_{H}$ or $g^{\prime}x_{\Phi}$ in the kinetic part of the Lagrangian. The other scenario $U(1)_{q+xu}$ reduces to B--L when $x=1$ and it can not reduce to the $U(1)_{R}$ case because $U(1)_{q+xu}$ charges of the SM right handed fermions do not become zero simultaneously.
The renormalizable scalar potential in this scenario can be written as
@@
where $q_{f_{L}}$ is the $U(1)_{X}$ charge of the SM lepton doublet, $Z^{\prime}$ gauge boson can decay into a pair of heavy Majorana neutrinos and the partial decay width is given by
@@ IV Results and discussions
$\frac{g_{X}^{2}}{{M_{Z^{\prime}}}^{2}-s}[\overline{e}\gamma^{\mu}(\tilde{x}_{\ell}P_{L}+\tilde{x}_{e}P_{R})e][\overline{f}\gamma_{\mu}(\tilde{x}_{f_{L}}P_{L}+\tilde{x}_{f_{R}}P_{R})f]\,,$ (56)
@@
$M_{Z^{\prime}}^{2}\ \gtrsim\ \frac{{g_{X}}^{2}}{4\pi}|{x_{e_{A}}}x_{f_{B}}|(\Lambda_{AB}^{f\pm})^{2}\,,$ (57)
@@
We calculate limits on $g_{X}-M_{Z^{\prime}}$ plane in the $U(1)_{X}$ scenario for different $x_{H}$ from the dilepton and dijet searches in ATLAS and CMS ATLAS:2019erb ; CMS:2019tbu experiments of the LHC. The corresponding cross sections for these scenarios are estimated as $\sigma_{\rm Model}$ involving the $Z^{\prime}$ contribution for different $x_{H}$ from $U(1)_{X}$ model considering a trial value of the general $U(1)_{X}$ gauge coupling $g_{\rm{Model}}$ at $13$ TeV varying $M_{Z^{\prime}}$. Comparing these estimated cross sections with the observed cross sections from the LHC $(\sigma_{\rm{Obs.}})$ to estimate limits on the gauge coupling for different $M_{Z^{\prime}}$ and $x_{H}$ following
$g_{X}\ =\ \sqrt{g_{\rm{Model}}^{2}\left(\frac{\sigma_{\rm{Obs.}}}{\sigma_{\rm{Model}}}\right)}.$ (58)
to estimate $95\%$ constraints on the $g_{X}-M_{Z^{\prime}}$. To estimate the cross sections of $Z^{\prime}$ production we used narrow width approximation (NWA) as
[…]
where $q(\bar{q})(x,Q)$ is the parton distribution functions of the (anti)quark and $\hat{s}=xys$ is the invariant mass squared of the colliding quark at the center of mass energy $\sqrt{s}$ considering up-type and down-type quarks. Due to NWA, the cross section of the colliding quarks to produce $Z^{\prime}$ boson is
$\hat{\sigma}=\frac{4\pi^{2}}{3}\frac{\Gamma(Z^{\prime}\to q\overline{q})}{M_{Z^{\prime}}}\delta(\hat{s}-M_{Z^{\prime}})$ (60)
with a factorization scale at $Q=M_{Z^{\prime}}$ and employing CTEQ6L Pumplin:2002vw as the parton distribution function. Partial decay widths of $Z^{\prime}$, used in this analysis, into a pair of charged fermions are calculated as
$\Gamma(Z^{\prime}\to\bar{f}f)$ $=N_{C}\frac{M_{Z^{\prime}}g_{X}^{2}}{24\pi}\left[\left(q_{f_{L}}^{2}+q_{f_{R}}^{2}\right)\left(1-\frac{m_{f}^{2}}{M_{Z^{\prime}}^{2}}\right)+6q_{f_{L}}q_{f_{R}}\frac{m_{f}^{2}}{M_{Z^{\prime}}^{2}}\right]\sqrt{1-4\frac{m_{f}^{2}}{M_{Z^{\prime}}^{2}}},$ (61)
where $m_{f}$ stands for the mass of the SM fermions, $N_{C}$ is the color factor which is $1(3)$ for the SM leptons (quarks) and $q_{f_{L}(f_{R})}$ are the $U(1)_{X}$ charges of the left (right) handed charged fermions of the model. That into a pair of light neutrinos $(\nu_{L})$ is calculated as
$\Gamma(Z^{\prime}\to\nu\nu)=\frac{M_{Z^{\prime}}g_{X}^{2}}{24\pi}q_{\nu_{L}}^{2}\leavevmode\nobreak\ ,$ (62)
neglecting the tiny mass of the light neutrinos. In this model, the $Z^{\prime}$ gauge boson can also decay into a pair of heavy Majorana neutrinos when the $Z^{\prime}$ is heavier than twice the mass of the heavy neutrinos $(m_{N})$ which is a free parameter. The corresponding partial decay width of $Z^{\prime}$ into single generation of heavy neutrino pair is given by
$\Gamma(Z^{\prime}\to N_{R}N_{R})=\frac{M_{Z^{\prime}}g_{X}^{2}}{24\pi}q_{N_{R}}^{2}\left(1-4\frac{m_{N}^{2}}{M_{Z^{\prime}}^{2}}\right)^{\frac{3}{2}}\leavevmode\nobreak\ ,$ (63)
with $q_{N_{R}}$ is the general $U(1)_{X}$ charge of the heavy neutrinos, and in our case, it is $q_{N_{R}}=-x_{\Phi}=-1$. If we consider $m_{N}>M_{Z^{\prime}}/2$, then $Z^{\prime}\to N_{R}N_{R}$ mode will be kinematically forbidden. We estimate $\sigma_{\rm{Model}}=$ $\sigma(pp\to Z^{\prime})\times$ BR$(Z^{\prime}\to 2\ell)$ for dilepton and $\sigma(pp\to Z^{\prime})\times$ BR$(Z^{\prime}\to 2j)$ for dijet final states, respectively. Total decay width of $Z^{\prime}$ used in this analysis has been calculated using the partial decay widths following Eqs. (61)- (63) to obtain the respective branching ratios. Finally bounds on $g_{X}-M_{Z^{\prime}}$ plane are calculated from Eq. (58) and shown in Fig. 2. The dilepton bounds provide strong constraints for different $x_{H}$ compared to the dijet bounds.
@@
Finally from Fig. 2 we compare the limits obtained from resonant leptogenesis and collider searches for different $x_{H}$. We find that limits obtained from resonant leptogenesis become stronger than the limits obtained from LEP-II, LHC for $M_{Z^{\prime}}>5.8$ TeV for $x_{H}=-2$ and $-1$. The constraints obtained from LEP-II become stronger than the CMS limits for $x_{H}\leq 0$, ruling out some parts of the resonant leptogenesis parameters for $x_{H}=0$ and $1$, respectively. We find that For $x_{H}=2$, LEP-II limits rule out the limits obtained from the resonant leptogenesis. We find that prospective limits in case of 250 GeV ILC can be comparable with the limits obtained by the resonant leptogenesis scenario. Hence these limits could be probed at 250 GeV ILC for $x_{H}=-1$. Other prospective limits from ILC at different center of mass energies more than 250 GeV and different $x_{H}$ could provide stronger constraints than the limits obtained from resonant leptogenesis.